OpenStudy (anonymous):

I need step by step help to this question: sqrt 363 -3 sqrt27

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

You would have to make the numbers inside the radicals the same, or find the LCM

OpenStudy (matthew120):

here baby

OpenStudy (matthew120):

https://mathway.com/popular-problems/Algebra/200675

OpenStudy (anonymous):

basically (121 x 3) is 363

OpenStudy (anonymous):

That means our LCM is 3

OpenStudy (anonymous):

(9 x 3) is 27

OpenStudy (anonymous):

\[\sqrt{11^{2 }\times3} - 3\sqrt{3^{2}\times3}\]

OpenStudy (anonymous):

You still here?

OpenStudy (anonymous):

Simplifying it would result in: \[11\sqrt{3} - \left( 3\times3 \right)\sqrt{3}\]

OpenStudy (matthew120):

barf bag dont get any ideas

OpenStudy (anonymous):

\[11\sqrt{3} - 9\sqrt{3} = 2\sqrt{3}\]

OpenStudy (anonymous):

done

OpenStudy (anonymous):

Ok @matthew120, calm down buddy. Im working out the problem for her

OpenStudy (anonymous):

@BarfBag Thanks :)

OpenStudy (matthew120):

lol jk man

OpenStudy (anonymous):

Np